SUB-DOPPLER AND FTMW SPECTROSCOPY OF HC_{3}N ISOTOPOLOGUES

\underline{A.V. LAPINOV}, G. YU. GOLUBIATNIKOV, Institute of Applied Physics of RAS, 46 Ulyanov str., 603950 Nizhny Novgorod, Russia; A. P. VELMUZHOV, Institute of Metalloorganic Chemistry of RAS, 49 Tropinin str., 603950 Nizhny Novgorod, Russia; J.-U. GRABOW, Institute of Physical Chemistry and Electrochemistry, Leibniz University of Hannover, Callinstrasse 3A, 30167 Hannover, Germany; and A. GUARNIERI, Technical Faculty of Christian Albrecht University of Kiel, Kaiserstrasse 2, 24143 Kiel, Germany.

We report results of precise sub-Doppler spectroscopy of HC_{3}N, H^{13}CCCN, HC^{13}CCN, HCC^{13}CN and HCCC^{15}N at 45−510 GHz with Lamb-dip spectrometer of IAP RAS. Hf structure of rotational transitions of all above species below 27 GHz as well as inversion transitions of NH_{3}(1,1) and (2,2) were measured using FTMW spectrometer of University of Hannover. New comparison of HC_{3}N and NH_{3} laboratory frequencies with radio astronomical observations of dark clouds^{a}^{,}^{b} shows an upper limit of m_{e}/m_{p} variation in our Galaxy as ≤ 3 · 10^{−9}.

aS. A. Levshakov, A. V. Lapinov, C. Henkel et al. Astron. Astrophys. \underline{524}, A32, 2010.

bS. A. Levshakov, P. Molaro, A. V. Lapinov et al. Astron. Astrophys. \underline{512}, A44, 2010.